OpenStudy (anonymous):

L1 passes through (1,4) and (3,-1) while L2 passes through (3,2) and (5,-1). find the angle from L1 to L2.

OpenStudy (zale101):

first, let us identify the slope so our equation will look like y=mx+b, we need M which is the slope

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

can you teach me step by step "?

OpenStudy (anonymous):

the slope of each line can be found by using the point-slope formula: \[y_2 - y_1 = m(x_2 - x_1)\] I'll show you for the first line, then you can do similarly for the second line on your own I take the points, (1,4) and (3,-1) and plug them in to the right spots in the equation -1-4 = m(3-1) -5 = m*2 -5/2 = m [here is the slope for the first line] you try it for the second line which has points, (3,2) and (5,-1).

OpenStudy (zale101):